Equivalent Units of Production

The Converting Department of Hopkinsville Company had 1,080 units in work in process at the beginning of the period, which were 60% complete. During the period, 22,800 units were completed and transferred to the Packing Department. There were 1,200 units in process at the end of the period, which were 60% complete. Direct materials are placed into the process at the beginning of production.

Determine the number of equivalent units of production with respect to direct materials and conversion costs. If an amount is zero, enter in "0".

Hopkinsville Company
Number of Equivalent Units of Production
Whole Units Direct Materials Equivalent Units Conversion Equivalent Units
Inventory in process, beginning
Started and completed
Transferred to Packing Department
Inventory in process, ending
Total
Answer #1
Hopkinsville Company
Number of Equivalent Units of Production
Whole Units Direct Materials Equivalent Units Conversion Equivalent Units
Inventory in process, beginning 1080 0 432
Started and completed 21720 21720 21720
Transferred to Packing Department 22800 21720 22152
Inventory in process, ending 1200 1200 720
Total 24000 22920 22872
Notes
Inventory in process, beginning: 1080 X 40% or 1080 X .40= 432   Started and completed: 22800-1080=21720 Transferred to Packing Department: 21720+1080=22800, 21720+0=21720,432+21720=22152 Inventory in process, ending: (60%), 1200,1200, 60% of 1200 or 1200 X .60=720